### v2

# LLM4MSR: An LLM-Enhanced Paradigm for Multi-Scenario Recommendation

Abstract.
As the demand for more personalized recommendation grows and a dramatic boom in commercial scenarios arises, the study on multi-scenario recommendation (MSR) has attracted much attention, which uses the data from all scenarios to simultaneously improve their recommendation performance. However, existing methods tend to integrate insufficient scenario knowledge and neglect learning personalized cross-scenario preferences, thus leading to sub-optimal performance. Meanwhile, though large language model (LLM) has shown great capability of reasoning and capturing semantic information, the high inference latency and high computation cost of tuning hinder its implementation in industrial recommender systems. To fill these gaps, we propose an LLM-enhanced paradigm LLM4MSR in this work. Specifically, we first leverage LLM to uncover multi-level knowledge from the designed scenario- and user-level prompt without fine-tuning the LLM, then adopt hierarchical meta networks to generate multi-level meta layers to explicitly improve the scenario-aware and personalized recommendation capability. Our experiments on KuaiSAR-small, KuaiSAR, and Amazon datasets validate significant advantages of LLM4MSR: (i) the effectiveness and compatibility with different multi-scenario backbone models, (ii) high efficiency and deployability on industrial recommender systems, and (iii) improved interpretability. The implemented code and data is available to ease reproduction [^1] https://github.com/mindspore-lab/models/tree/master/research/huawei-noah/LLM4MSR, [^2] https://github.com/Applied-Machine-Learning-Lab/LLM4MSR.
Large Language Model, Multi-Domain, Multi-Scenario Recommendation, Click-Through Rate Prediction
🖂Corresponding author
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679743isbn: 979-8-4007-0436-9/24/10ccs: Information systems Recommender systems
1. Introduction
The research on recommender system (RS) has been a hot spot in the past few years which targets at mining users’ interests from the vast amount of historical interaction data. In order to provide services in a more personalized manner and make profits, an explosive growth of business scenarios (also known as domains) are emerging in commercial RSs such as the search and recommendation services on mobile apps (Gong et al., 2023; Sun et al., 2023) and different product categories on the e-commerce platform (Ni et al., 2019). Therefore, as a group of efficient joint modeling approach, multi-scenario recommendation (MSR) has drawn attention increasingly, which exploits all data to improve the recommendation accuracy on all these scenarios, thus addressing the data sparsity issue (Zhu et al., 2021) and reducing computation cost.
Nevertheless, existing MSR methods usually suffer from the following two limitations: (i) Insufficient scenario knowledge is incorporated. Specifically, domain indicator is usually the only scenario knowledge leveraged to capture the domain distinction in practice (Sheng et al., 2021; Chang et al., 2023), while the abundant semantic information such as the description of scenario is neglected. Consequently, the correlations among scenarios are not adequately established. (ii) Users’ personalized preferences across scenarios tend to be ignored, since most MSR models merely rely on different parameter sharing patterns in multi-task learning (Sheng et al., 2021; Wang et al., 2023a; Liu et al., 2023b) and the collaborative signal learned in conventional RSs to conduct recommendation. However, these deficiencies are still overlooked even if they would lead to the sub-optimal performance of multi-scenario recommender system.
To tackle these drawbacks of conventional MSR models, we resort to large language model (LLM) which is known for its remarkable capability of language understanding and reasoning (Fan et al., 2023). Although there have been some attempts to enhance RSs harnessing LLM (Gong et al., 2023; Yin et al., 2023; Li et al., 2023a), there are some key challenges. The first is high inference latency of LLM since the existing methods conduct training and inference with LLM through sample level alignment (Li et al., 2023a; Wang et al., 2023b). Second, the information generated from LLM is usually in natural language format. In other words, it requires filling the gap between semantic and recommendation space like employing additional knowledge encoder (Xi et al., 2023) and instruction tuning on LLM (Yin et al., 2023) which brings high computation cost. Moreover, both conventional RS and LLM-based recommendation methods lack interpretability.
To this end, we propose an LLM-based paradigm to enhance conventional multi-scenario recommendation, which leverages LLM and hierarchical meta networks to explicitly improve the performance of the backbone model on all scenarios. To be specific, on the one hand, LLM helps grasping the cross-scenario correlation and personalized preferences, while the meta networks act as a flexible and adaptive bridge connecting the semantic space in LLM and the recommendation space in the multi-scenario backbone model. On the other hand, to tackle the efficiency of LLM (see Section 3.4), a frozen LLM is used without tuning and high level of user- & scenario knowledge is reasoned about through prompt design.
The main contributions of this paper are summarized as follows:
  • We propose an effective efficient interpretable paradigm LLM4MSR for multi-scenario recommendation enhancement by exploiting LLM. To the best of our knowledge, it is the first practical solution to reasoning about multi-scenario knowledge through LLM and bring information gain in recommender systems.
  • The scenario commonality and heterogeneity, and users’ cross-scenario preferences are explicitly captured by LLM through the designed scenario- and user-level prompt. Furthermore, they are leveraged by the hierarchical meta networks which generate meta layers to explicitly enhance the scenario-aware and personalized recommendation capability of multi-scenario backbones.
  • Extensive experiments on three real-world datasets show that LLM4MSR is an effective paradigm compared with state-of-the-art multi-scenario enhancement methods and compatible with various MSR backbone models. Besides, it is efficient to deploy on industrial recommender systems and enables real-time recommendation without fine-tuning LLM.
2. Preliminary
This section first introduces the problem formulation of multi-scenario CTR prediction and the typical architecture of multi-scenario backbone models. Then, large language model is illustrated.
2.1. Multi-Scenario CTR Prediction
The Click-Trough Rate (CTR) prediction task is essential in recommendation which is a binary classification problem. Specifically, in a multi-scenario recommender system, the historical interaction sample $(d,\bm{x})$ is taken as input to predict the ground truth label $y$ where $y=1$ and $y=0$ denote click and unclick, respectively. $d$ represents the domain indicator $d\in\{1,2,\dots,D\}$ indicating the origin of sample. $\bm{x}$ denotes the feature fields including user and item attributes. For simplicity, suppose there are $M$ categorical features, then $\bm{x}=[\bm{x}_{1},\ldots,\bm{x}_{m},\ldots,\bm{x}_{M}]$ where $\bm{x}_{m}$ indicates the one-hot encoding of the $m$-th feature. Next, $\bm{x}$ is mapped into a dense vector $\bm{e}=[\bm{e}_{1}\|\ldots\|\bm{e}_{m}\|\ldots\|\bm{e}_{M}]$ through an embedding layer where $\|$ denotes the concatenation operation. For the $m$-th feature, we calculate $\bm{e}_{m}$ through a look-up operation $\bm{e}_{m}=\bm{E}_{m}\cdot\bm{x}_{m}$ in which $\bm{E}_{m}\in\mathbb{R}^{u_{m}\times Dim}$ denotes the embedding table, $u_{m}$ denotes the value counts of feature, and $Dim$ is the embedding size. Finally, the prediction result $\hat{y}$ is obtained via $\hat{y}=f_{d}(\bm{e})$ where $f_{d}$ represents the recommendation model applied to the $d$-th scenario. The loss function is binary cross entropy loss or the so-called Logloss (Berkson, 1944):
(1) $\underset{\Theta}{\min}\mathcal{L}=-\frac{1}{B}\sum_{i=1}^{B}y_{i}\log\left(\hat{y}_{i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{i}\right),$
where $\Theta$ is the set of parameters learned, $B$ represents the number of batch samples, and $y_{i}$ and $\hat{y}_{i}$ denote the true label and prediction of the $i$-th sample, respectively.
As depicted in Figure 1, most multi-scenario recommendation models model the commonality and distinction among scenarios through explicitly maintaining domain-shared parameters $\Theta_{s}$ like the experts in MMoE (Ma et al., 2018) and domain-specific parameters $\Theta_{u}$ like towers and gates in PLE (Tang et al., 2020).
Figure 1. Demonstration of multi-scenario recommender system. The blue part represents shared parameters while red and yellow denote scenario-specific parameters.
3. Proposed Method
Figure 2. Overall framework of LLM4MSR. Different colors of solid arrows indicate different data flow. In multi-scenario knowledge reasoning step, an LLM captures scenario correlation and cross-scenario preferences from scenario- and user-level prompt without fine-tuning. In multi-level knowledge fusion step, two hierarchical meta networks, which adaptively generate scenario- and user-level meta layers in parallel and at bottom, are end-to-end trained with the multi-scenario backbone model.
We first illustrate the overall framework of LLM4MSR. Next, it is further detailed in the multi-scenario knowledge reasoning and multi-level knowledge fusion step. Finally, we discuss the idea and characteristics of LLM4MSR compared with the existing multi-scenario and LLM-based recommendation methods, and illustrate the necessity of LLMs versus Pre-trained Language Models (PLMs).
3.1. Overall Framework
To tackle the limitations of the existing MSR methods, i.e., insufficient scenario knowledge incorporation and ignoring personalized cross-scenario interest, we propose a paradigm LLM4MSR for multi-scenario enhancement which is compatible with different multi-scenario backbone models. Specifically, it is equipped with an LLM with frozen parameters and two learnable meta networks. Meanwhile, the general procedure of LLM4MSR is summarized in Algorithm 1. It is comprised of two steps: (i) multi-scenario knowledge reasoning (from Line 1 to 10) including scenario correlation and cross-scenario interests inference, and (ii) multi-level knowledge fusion (from Line 11 to 15) to adaptively fuse with backbone.
1
2
Input: Domain ID $d\in\{1,2,\dots,D\}$; set of user $u\in\{1,2,\dots,U\}$; features $\bm{x}$; true label $y$; a pre-trained large language model (LLM)
Output: A well trained model for all scenarios
3
Step 1: Multi-scenario Knowledge Reasoning
4for $d\in\{1,2,\dots,D\}$ do
5   Generate scenario-level prompt;
6   Feed into LLM;
7   Obtain the last hidden state of scenario-level output;
8
9 end for
10
11for $u\in\{1,2,\dots,U\}$ do
12   Generate user-level prompt;
13   Feed into LLM;
14   Obtain the last hidden state of user-level output;
15
16 end for
17
Step 2: Multi-level Knowledge Fusion
18while not converge do
19   Sample a mini-batch data from all scenarios;
20
21  Calculate the corresponding scenario- and user-level meta layers via Equation (3) and (4);
22   Calculate the prediction result via Equation (7);
23   Calculate the loss via Equation (1);
24   Take the gradient and update parameters;
25
26 end while
Algorithm 1 LLM4MSR paradigm taking multi-scenario CTR prediction as an example
The illustration of LLM4MSR is depicted in Figure 2. Specifically, the scenario-level prompt is generated from the scenario statistics, semantic descriptions, and expert knowledge while the user-level prompt is constructed using historical interactions. Afterward, they are fed into LLM to generate scenario correlation and personalized cross-scenario interest respectively. Notably, only the last hidden state of LLM output (i.e., the hidden state of the last token generated) is chosen as representative and stored which is a high-dimensional dense vector, i.e., in 4096 dimensions taking ChatGLM2-6B (Du et al., 2022) as an example. It is common because the generation of the last token is based on all previous ones and contains overall information. Finally, the hierarchical meta networks produce meta layers in scenario- and user-level, and they are trained with the MSR backbone model to enhance its all-scenario performance. Overall, the final enhanced model can be obtained in a simple yet efficient end-to-end training manner by optimizing Equation (1).
3.2. Multi-Scenario Knowledge Reasoning
On the one hand, it is difficult for the mainstream MSR models to grasp the complex scenarios correlations because in practice the domain indicator distinguishing data origin is usually the only scenario-related knowledge (Sheng et al., 2021; Chang et al., 2023; Gao et al., 2023), thus their scenario representation is only learned from the implicit collaborative signals. Therefore, we design a scenario-level prompt to explicitly capture the commonality and distinction of scenarios from their statistics and semantic descriptions. For example, as shown in the purple box in Figure 2, the statistics like the number of interactions, users, items, and overlaps on KuaiSAR-small together with the background semantic information are given to uncover both similarity and heterogeneity of the recommendation and search scenario.
On the other hand, existing MSR models tend to neglect modeling personalized interest and it is challenging to capture user’s cross-scenario preferences in a disentangled manner. To this end, we design a user-level prompt to explicitly explore personalized common and distinct interests among all the scenarios from the abundant historical interaction. Intuitively, positive interaction behavior tend to better reflect user’s interest than negative behavior. Meanwhile, the proportion of negative samples is usually large and they contain much noise information due to the biases and uncertainty in data collection and user interaction. Consequently, the positive interaction data (i.e., samples with positive interaction behavior such as click, forward, and like) are filtered and leveraged in the user-level prompt. As shown in the blue box in Figure 2(a), LLM helps explicitly capturing user’s preferences in the search and recommendation scenario considering both the interaction frequency and the semantic relationships, e.g., in the category and title of item. The specific prompt template are provided in Appendix B.
After obtaining the scenario- and user-level prompt, they are passed into LLM to reason about the multi-scenario knowledge. Moreover, we conduct a case study in Section 4.6 to investigate what multi-scenario knowledge can LLM obtain and how it contributes to enhancing the MSR backbone model.
3.3. Multi-Level Knowledge Fusion
Existing methods either choose to directly leverage the augmented information from LLM as additional input features in the RS (Xi et al., 2023), adopt a linear projection or a fully connected network to conduct dimension transformation and align instance representation (Li et al., 2023a), or fine-tune the LLM (Zheng et al., 2023). However, we argue that these are sub-optimal solutions (Su et al., 2023). First, since the LLM-augmented information is treated as a common feature by recommender system, it is unable to play a significant role as expected because the number of feature fields is usually large (Zhao et al., 2021). Second, simply aligning the dimension would result in loss of information (Zebari et al., 2020). Third, tuning LLM brings high cost (Chen et al., 2023b, a) and it requires consistent update.
Therefore, after obtaining the informative multi-scenario knowledge which is in the form of high-dimensional hidden vectors from LLM, our proposed LLM4MSR adopts a new and efficient approach leveraging hierarchical meta networks to bring information gain. To be specific, the meta networks dynamically generate meta layers to adaptively fuse the scenario- and user-level knowledge with the collaborative information in the MSR backbone model. Intuitively, on the one hand, the scenario-level knowledge usually plays a more general role like the domain distinction captured by auxiliary network in STAR (Sheng et al., 2021). On the other hand, the user-level information tends to be more personalized and significant (Chang et al., 2023) because it breaks the boundary of scenarios and contains a complete user profile across all scenarios. Consequently, we propose a hierarchical ‘bottom + parallel’ structure, in which the user-level meta layers are placed at the bottom while the scenario-level meta layers are implemented in parallel to the MSR backbone, as shown in Figure 2(b). Notably, They target at improving scenario-aware recommendation and cross-scenario personalization capability respectively, thus contributing to knowledge fusion and enhancement in different levels and granularity. We will empirically investigate this architecture of LLM4MSR through experiments in Section 4.3 where we also observe that user-level meta layers at bottom and scenario-level meta layers in parallel is the optimal combination.
Specifically, denote the last hidden state of LLM output as $\bm{h}_{LLM}$, the meta work takes it as input and outputs $\bm{h}_{mw}\in\mathbb{R}^{Dim_{mw}}$ and $\bm{h}_{mb}\in\mathbb{R}^{Dim_{mb}}$, then a reshape operation is employed to obtain $K$ meta layers in total:
(2) $\bm{h}_{mw},\bm{h}_{mb}=\operatorname{Meta~Network}(\bm{h}_{LLM}),$
(3) $\bm{W}_{l}^{(i)}=Reshape(\bm{h}_{mw})$
(4) $\bm{b}_{l}^{(i)}=Reshape(\bm{h}_{mb}),i\in\{1,2,\ldots,K\}$
[…]
(5) $\bm{h}^{(i)}=\sigma(\bm{W}_{l}^{(i)}\bm{h}^{(i-1)}+\bm{b}_{l}^{(i)}),i\in\{1,2,\ldots,K\}$
where $\bm{h}^{(i)}$ is the output of the $i$-th meta layer. Afterward, the feature embedding $\bm{e}$, which is exactly $\bm{h}^{(0)}$, is transferred into the obtained meta layers. Finally, the prediction result is calculated by combining the output of the parallel scenario-level meta layers and the result from MSR backbone in an adaptive manner:
(6) $\bm{h}=\operatorname{MSR}(\bm{h}_{u}^{(K)}),$
(7) $\hat{y}=\sigma^{\prime}(\alpha\cdot\bm{h}_{s}^{(K)}+(1-\alpha)\cdot\bm{h}),$
where $\bm{h}$ is the output hidden vector of the MSR backbone model, $\bm{h}_{u}^{(K)}$ is the output of the $K$-th user-level meta layer, $\bm{h}_{s}^{(K)}$ is the output of the $K$-th scenario-level meta layer, $\sigma^{\prime}$ is the activation function for the final task (e.g., sigmoid function for CTR prediction task), and $\alpha$ is a learnable parameter.
3.4. Discussion
In the following part, we first illustrate the idea of the proposed LLM4MSR paradigm, then compare it with the mainstream multi-scenario RSs and the LLM-based recommendation methods, and finally demonstrate the necessity of LLM here compared with PLM.
Basically, rather than designing a specific multi-scenario recommendation model architecture, our goal is to propose a general multi-scenario enhancement paradigm. Specifically, considering the model architecture and optimization, LLM4MSR does not impose any restrictions on the multi-scenario backbone models, thus being model-agnostic. Meanwhile, for LLM the most two significant questions are: (i) WHAT information can be obtained from LLM and (ii) HOW to use it to bring information gain, which correspond to the multi-scenario knowledge reasoning and multi-level knowledge fusion steps proposed in our LLM4MSR.
LLM4MSR v.s. multi-scenario RS. Conventional multi-scenario RSs treat the domain indicator either as a distinct feature (Sheng et al., 2021) or as the only scenario knowledge to generate scenario-specific module (Gao et al., 2023), and adopt different parameter sharing patterns to model scenario correlations where only the collaborative signal is learned. By contrast, LLM4MSR generates LLM-enhanced scenario- and user-level representation from the semantic space, then it dynamically generates meta layers to effectively and efficiently combine with the multi-scenario RS based on these representations. Consequently, the LLM-augmented information is flexibly fused with the collaborative signal in conventional RS, thus achieving enhancement.
LLM4MSR v.s. LLM-based RS. Existing LLM-based recommender systems can be categorized into two groups, i.e., using LLM as the knowledge enhancer (Li et al., 2023a; Wang et al., 2023b; Lin et al., 2023; Gong et al., 2023) and directly as the recommender (Geng et al., 2022; Cui et al., 2022; Liu et al., 2023a). However, the former models possess low efficiency since they require aligning the information from LLM and RS in the instance level. Meanwhile, the latter methods lack scenario knowledge, thus achieving unsatisfying recommendation performance and suffering from high inference latency of LLM.
By contrast, LLM4MSR can efficiently and effectively tackle these problems. First, thanks to the zero-shot capability (Wang et al., 2022b), an LLM with fixed parameters is leveraged free from tuning in (Yin et al., 2023; Li et al., 2023a; Wang et al., 2023b). Meanwhile, the way of obtaining information from LLM is shifted from the low ‘instance-level’ (e.g., in (Li et al., 2023a; Wang et al., 2023b; Zheng et al., 2023)) to the high ‘user- & scenario-level’, thus improving efficiency. Besides, through hierarchical meta networks, the multi-level scenario knowledge is adaptively injected and lead to enhancement. Moreover, apart from distinct scenario-specific and user-specific features, it also enables to incorporate expert scenario knowledge through the prompt design, which is more universal and advanced than simply extracting invariant features and aligning the heterogeneous ones in (Gong et al., 2023).
Necessity of LLMs v.s. PLMs. Unlike PLMs like BERT trained with masked language modeling task, most popular LLMs like ChatGLM2 (Du et al., 2022) are trained with next token prediction objective. Therefore, rather than simply deriving a representation of existing (old) information by adopting PLMs as encoder (like BERT in CTRL (Li et al., 2023a)), the key of leveraging LLM as reasoner in LLM4MSR is to further infer useful (new) knowledge like analyzing user’s cross-scenario preference. This knowledge (in the form of hidden vectors) could potentially help enhancing conventional recommender systems. Besides, a case study is conducted in Section 4.6 and the knowledge (in the form of text) from ChatGLM2-6B is shown in Figure 5, which helps improving interpretability.
By contrast, we also adopt GPT-2 (a typical PLM) to reason about user_1’s interest given the same input prompt as in ChatGLM2-6B, and it generates the following nonsense text:
"The main strategy of this study was to combine two
scenarios into a two-way analysis. The primary strategies
are presented in Table I. If the main scenario is presented
as a list of stories, then..."
In that case, such information could not bring information gain and improve the recommendation performance.
4. Experiments
Extensive experiments are conducted on three public datasets to verify the effectiveness and efficiency of our proposed LLM4MSR paradigm, and the following five questions are answered:
  • RQ1: Is LLM4MSR an effective paradigm and compatible with different multi-scenario backbone models?
  • RQ2: What is the effect of the scenario-level and user-level prompt? What is their optimal architecture correspondingly?
  • RQ3: What is the impact of interaction threshold and the number of neural network layers generated by the meta network?
  • RQ4: What is the efficiency of LLM4MSR compared with the original multi-scenario backbone models?
  • RQ5: How does LLM help improving the multi-scenario backbone models through our proposed LLM4MSR?
4.1. Experimental Settings
4.1.1. Datasets
Our experiments are conducted on three datasets: KuaiSAR-small, KuaiSAR (Sun et al., 2023), and Amazon (Ni et al., 2019). To be specific, there is a search and a recommendation scenario on KuaiSAR-small and KuaiSAR while there are three correlated recommendation scenarios selected on Amazon. Furthermore, their statistics and detailed descriptions are summarized in Table 1. To be specific, we divide the original data into training, validation, and test set with the ratio of 8:1:1 by chronological order. Besides, the sparsity metric denotes the proportion of unclicked samples with label $y=0$.
  • KuaiSAR-small & KuaiSAR [^3] https://kuaisar.github.io/ datasets are both crawled from the short-video platform Kuaishou. The only difference is KuaiSAR-small is a smaller version of data in 9 days while KuaiSAR is in 19 days. Specifically, all users are overlapped in the recommendation and search scenario. Meanwhile, there are 97981 and 181849 items overlapped on KuaiSAR-small and KuaiSAR, respectively.
  • Amazon [^4] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/ is collected from Amazon. Following (Cui et al., 2020), we select three connected scenarios: All Beauty, Clothing Shoes and Jewelry, and Luxury Beauty and we also filter the overlapped users among these scenarios without overlapped items. The goal is to predict whether a user would give a rating (from 1 to 5) higher than 3 to an item.
Table 1. The statistics of KuaiSAR-small, KuaiSAR, and Amazon, where Rec denotes recommendation.
Dataset KuaiSAR-small KuaiSAR Amazon
Scenarios Rec Search Rec Search Rec#1\#1 Rec#2\#2 Rec#3\#3
Users 25,877 25,877 25,877 25,877 24,752 24,752 24,752
Items 2,281,034 1,974,165 4,046,363 2,974,596 8,788 193,304 6,980
Interactions 7,493,101 3,038,362 14,605,712 4,828,690 33,929 370,840 56,356
Sparsity 50.49% 87.87% 50.40% 87.88% 21.50% 19.61% 20.67%
4.1.2. Evaluation Metrics
Area Under the ROC curve (AUC) is adopted to conduct evaluation on the test set, where a higher AUC value at 0.001 level usually denotes significant enhancement (Cheng et al., 2016; Wang et al., 2023c).
4.1.3. Backbone Models
We implement LLM4MSR on the following representative multi-scenario backbone models commonly deployed in industrial recommender system:
  • STAR (Sheng et al., 2021) proposes a star topology architecture with a shared centered network multiplied by many domain-specific networks.
  • OMoE (Jacobs et al., 1991) leverages a group of expert networks and a gating network learned to ensemble the results from these experts.
  • MMoE (Ma et al., 2018) explicitly learns a gating network for each task to assemble task-specific information compared with OMoE.
  • PLE (Tang et al., 2020) proposes Customized Gate Control (CGC) which separates shared and specific expert networks and employs multi-level extraction network with a progressive routing mechanism.
  • AITM (Xi et al., 2021) adaptively learns to transfer information for different tasks with sequential dependence.
  • Shared Bottom is a multi-task model where the embedding layer is usually shared on which task-specific networks are built.
Table 2. Overall performance comparison of AUC on KuaiSAR-small, KuaiSAR, and Amazon dataset. The suffix ‘DN’, ‘EP’, and ‘ours’ denote Dynamic Network, EPNet, and LLM4MSR. Boldface denotes the highest value while underline indicates the second best result. $\star$ represents statistical significance with $p$-value <0.05<0.05 in $t$-test compared with the backbone model.
Backbones KuaiSAR-small KuaiSAR Amazon
Rec Search Rec Search Rec#1\#1 Rec#2\#2 Rec#3\#3
STAR 0.7225 0.6089 0.7387 0.6268 0.6156 0.6320 0.6225
[…]
STAR_ours $0.7276^{\star}$ $0.6181^{\star}$ $0.7408^{\star}$ $0.6332^{\star}$ $0.8720^{\star}$ $0.6364^{\star}$ $0.7543^{\star}$
OMoE 0.7241 0.6163 0.7394 0.6310 0.5995 0.6043 0.6252
[…]
OMoE_ours $0.7265^{\star}$ $0.6186^{\star}$ $0.7413^{\star}$ $0.6332^{\star}$ $0.8182^{\star}$ $0.6180^{\star}$ $0.6823^{\star}$
MMoE 0.7235 0.6150 0.7392 0.6313 0.5907 0.5999 0.6032
[…]
PLE 0.7249 0.6149 0.7396 0.6313 0.6059 0.6127 0.5995
[…]
PLE_ours $0.7269^{\star}$ $0.6184^{\star}$ $0.7408^{\star}$ $0.6343^{\star}$ $0.8265^{\star}$ $0.6250^{\star}$ $0.7074^{\star}$
AITM 0.7236 0.6154 0.7398 0.6329 0.6039 0.6126 0.6046
[…]
AITM_ours $0.7273^{\star}$ $0.6178^{\star}$ $0.7407^{\star}$ $0.6349^{\star}$ $0.8196^{\star}$ $0.6178^{\star}$ $0.7089^{\star}$
Shared Bottom 0.7228 0.6169 0.7389 0.6321 0.6073 0.6077 0.6268
[…]
Shared Bottom_ours $0.7269^{\star}$ $0.6182^{\star}$ $0.7400^{\star}$ $0.6341^{\star}$ $0.8323^{\star}$ $0.6262^{\star}$ $0.7182^{\star}$
4.1.4. Baselines
There are few works aimed at explicitly enhancing multi-scenario recommendation. Besides, previous LLM-based models are not applicable in the same setting (refer to Section 5.2), e.g., LLM as recommender fails to meet real-time inference requirement. Consequently, we compare LLM4MSR with two typical paradigm as baseline methods:
  • Dynamic Network. It targets at adaptively generating model parameters in a fine-grained manner (Yan et al., 2022). In multi-scenario modeling (Zhang et al., 2022), the scenario knowledge is employed to generate scenario-specific attention and tower module.
  • EPNet. It is a gate structure proposed in PEPNet (Chang et al., 2023). Specifically, it takes domain-side information as input and produce scenario-specific gates to transform embedding.
4.1.5. Implementation Details
The implementation details of the experiments conducted are summarized in Appendix A and the code is available to ease reproduction1. Specifically, for the open-source LLM we choose ChatGLM2-6B [^5] https://huggingface.co/THUDM/chatglm2-6b (Du et al., 2022) as an example.
4.2. Overall Performance (RQ1)
To answer RQ1, we verify LLM4MSR’s compatibility with various multi-scenario backbone models introduced in Section 4.1.3. Meanwhile, LLM4MSR is compared with two methods for multi-scenario recommendation enhancement as the baseline methods detailed in Section 4.1.4. The overall performance of AUC on KuaiSAR-small, KuaiSAR, and Amazon dataset is shown in Table 2.
First, Dynamic Network and EPNet are able to simultaneously improve the recommendation accuracy on different scenarios in most cases, while there are exceptions like ‘AITM_DN’ on KuaiSAR and ‘STAR_EP’ on all datasets. By contrast, the proposed LLM4MSR significantly enhances the recommendation performance in all scenarios on the three datasets especially in Rec#1 and Rec#3 on Amazon. It also consistently outperforms the baseline methods except in the search scenario on KuaiSAR-small based on MMoE. The reason for its superiority is that, on the one hand, both the multi-scenario backbone models and baseline methods incorporate insufficient scenario knowledge, i.e., domain id only, and lack personalized modeling. On the other hand, LLM4MSR first utilizes LLM to explore domain correlation and infer personalized cross-scenario interest, then integrates meta network to explicitly enhance multi-scenario modeling capability. Besides, we observe that compared with the significant enhancement on scenario Rec#1 and Rec#3, LLM4MSR results in relatively subtle improvement (around 0.7%) on scenario Rec#2 on Amazon. The reason is that Rec#1 and Rec#3 are relatively sparse where each user has 1.4 and 2.3 interactions on average as shown in Table 1. Therefore, less information of user preference on these two scenarios could be captured by LLM to enhance Rec#2.
Overall, LLM4MSR achieves an increase of 1.5%, 1%, and 40% in AUC on three datasets. To summarize, we can draw the conclusion that LLM4MSR serves as an effective and powerful enhancement paradigm, which possesses great compatibility and deployability on the mainstream multi-scenario recommendation backbone models.
4.3. Ablation & Component Study (RQ2)
To answer RQ2, we conduct ablation study and component analysis based on STAR considering the following variants.
  • user-bottom / user-parallel: It only adopts user-level meta layers at bottom / in parallel without scenario knowledge as input.
  • domain-bottom / domain-parallel: It only adopts scenario-level meta layers at bottom / in parallel without user knowledge.
  • LLM4MSR: Full model with both user-bottom and domain-parallel.
The results are depicted in Figure 3. On the one hand, the full model LLM4MSR achieves the best result on all scenarios on all the three datasets. The reason is that it makes full use of both the informative multi-scenario knowledge and the collaborative signal learned from the LLM and MSR backbone model, respectively. Meanwhile, from ‘user - bottom’ and ‘domain - parallel’ it is seen that both user- and domain-level prompt contribute to the performance enhancement on the backbone model. Nonetheless, the domain-level prompt plays a more significant role on KuaiSAR-small and KuaiSAR dataset, while the user-level prompt is a more essential component on Amazon dataset.
Figure 3. Ablation and component study on KuaiSAR-small, KuaiSAR, and Amazon dataset. The prefix ‘user’ and ‘domain’ denotes user-level and domain-level prompt while the suffix ‘bottom’ and ‘parallel’ indicates the architecture applied. We adopt the ratio of improvement in AUC value compared with the original STAR model as the backbone.
Figure 4. Analysis on the positive interaction threshold and meta layers evaluated by AUC on KuaiSAR-small dataset.
On the other hand, comparing ‘user - bottom’ and ‘user - parallel’ we can observe that the former is better while the latter structure even has a negative impact in many cases. Similarly, placing the layers generated from the domain-level prompt in parallel to the scenario tower of the backbone model is the better choice contrasting ‘domain - bottom’ and ‘domain - parallel’. Consequently, we can draw the conclusion that the combination of layers from user-level prompt at bottom and layers from domain-level prompt in parallel is the optimal architecture. A reasonable explanation is that the user-level multi-scenario information helps personalizing and modulating the general embedding better in a lower level. By contrast, the scenario-level information acts better in a higher level like the auxiliary network adopted in STAR (Sheng et al., 2021). Last but not least, it is intuitive to investigate the effect of LLM by replacing the obtained knowledge in the form of hidden states with randomly initialized vectors. However, such enhancement is not observed.
4.4. Hyper-parameter Analysis (RQ3)
We analyze the effect of two significant hyper-parameters, i.e., the threshold ($T$) of positive interactions in the user-level prompt and the number of fully connected layers ($K$) generated by the meta network using STAR as the backbone model on KuaiSAR-small dataset. Specifically, as shown in Figure 4(a), the recommendation accuracy continuously increases as $T$ rises until the token limit of LLM is reached, since more information about cross-scenario preference of users is introduced and captured by LLM. Besides, from Figure 4(b) we can see that the best performance is obtained where $K=2$. Nevertheless, the results degrade as the network becomes deeper and a possible reason is over-fitting.
4.5. Efficiency Analysis (RQ4)
Table 3. Efficiency comparison on KuaiSAR-small in seconds.
Models STAR STAR_DN STAR_EP STAR_ours
Training 2.8e-2 2.9e-2 2.9e-2 1.5e-1
Inference 1.7e-2 1.7e-2 1.7e-2 5.2e-2
To validate the efficiency of our proposed LLM4MSR paradigm, we take STAR as the backbone model and experiment on KuaiSAR-small dataset. As shown in Table 3, as an LLM-based method, LLM4MSR still enables real-time inference even if it requires more training and inference time than other baseline methods. Besides, the real-world industrial system has a maximum latency specification (let’s say 0.1s), and the present inference time (0.05s in our method) will not undermine the effectiveness of system considering both latency and success rate, while significant improvement of performance is still achieved. Consequently, LLM4MSR possesses an excellent trade-off in effectiveness and efficiency.
In practice, the last hidden states generated from the user-level prompt can be stored before training, since inferring user preference is the most time-consuming step in LLM4MSR and it takes about 2 to 4 seconds per user on one single GPU. Besides, the architecture with scenario-level parameters only (i.e., ‘domain - parallel’ in Section 4.3) can serve as an alternative approach because the scenario-level prompt does not require frequent updates or personalization, thus leading to higher efficiency.
4.6. Case Study (RQ5)
Figure 5. Case study of user_1 on KuaiSAR-small dataset. The purple and blue box denote the result on the scenario correlation and the inferred cross-scenario preferences from scenario- and user-level prompt by ChatGLM2-6B.
To answer RQ5, we conduct a case study on KuaiSAR-small dataset taking user_1 as an example and the results generated from the corresponding designed scenario- and user-level prompt by ChatGLM2-6B are shown in Figure 5. Specifically, we can see that ChatGLM2-6B not only explicitly captures the commonality and distinction between the recommendation and search scenario, but it also clearly infers user_1’s common and distinct interests among these two scenarios with an accurate user profile summary. Afterward, the generated information is integrated hierarchically by the meta network and meta layers. Consequently, the performance of the multi-scenario backbone model is enhanced. Meanwhile, this knowledge in different levels from LLM also helps improving the interpretability of the original recommender system.
5. Related Work
This section brief summarizes the recent work on multi-scenario recommendation and adopting language model for CTR prediction.
5.1. Multi-Scenario Recommendation
Multi-scenario recommendation (Wang et al., 2023c, 2022a, 2024; Li et al., 2023c), also known as multi-domain or multi-target cross-domain recommendation (Zhu et al., 2021), uses the data from all scenarios to simultaneously improve their recommendation performance and tackle data sparsity problem. For example, STAR (Sheng et al., 2021) adopts a star-like framework with a shared centered network and scenario-specific networks similar to the parameter sharing pattern design in multi-task modeling (Wang et al., 2023a). M2M (Zhang et al., 2022) proposes a meta learning mechanism with hierarchically organized meta attention and meta tower module to grasp inter-scenario correlations and scenario-specific representation. In addition, to address the imperfectly double seesaw phenomenon, PEPNet (Chang et al., 2023) proposes embedding and parameter personalized network to dynamically scale the parameters based on the designed gate mechanism in a domain-specific and task-specific manner, respectively.
Facing the drawbacks of the aforementioned methods introduced in Section 1, i.e., insufficient integration of scenario knowledge and neglecting personalized interest across different scenarios, our proposed framework leverages LLM to uncover domain commonality and distinction and infer users’ cross-scenario preference, which are further utilized to enhance the multi-scenario backbone model.
5.2. Language Model for CTR Prediction
Several works try to incorporate Language Models (LMs) including Large Language Models (LLMs) and Pre-trained Language Models (PLMs) with recommendation for Click-Through Rate (CTR) prediction (Li et al., 2023a; Fu et al., 2023; Xi et al., 2023; Lin et al., 2023; Wang et al., 2023b; Gong et al., 2023) since they have the potential to bring information gain (Zhao et al., 2024; Xu et al., 2023; Li et al., 2023b; Liu et al., 2024). Specifically, one of the solutions is to treat the semantic knowledge as cross-modal information and align with the collaborative signal in conventional recommender systems. For example, CTRL (Li et al., 2023a) employs contrastive learning for knowledge alignment and integration. KAR (Xi et al., 2023) first generates reasoning knowledge of user preferences and item factual information, then encodes and adapts them to the backbone recommendation models. Besides, S&R Multi-domain Foundation Model (Gong et al., 2023) simply adopts LLM to extract scenario-invariant text features and tackle item heterogeneity issue.
However, on the one hand, these methods necessitate conducting both training and inference with LM in the sample level, leading to low efficiency of inference and poor deployability on industrial recommender systems. On the other hand, the connection between scenarios and personalized modeling are ignored. In this paper, we investigate a new LLM-enhanced paradigm, which not only improves the performance of multi-scenario modeling through the inferred scenario-level correlation and user-level cross-scenario interest, but it also enables efficient training and real-time inference.
6. Conclusion
In this paper, we propose an effective efficient interpretable paradigm LLM4MSR which first leverages LLM to reason about multi-scenario knowledge, including scenario commonality and distinction, and users’ cross-scenario preferences through the designed scenario- and user-level prompt. Then it adopts hierarchical meta networks which generate meta layers to adaptively fuse the multi-level knowledge and improve the conventional multi-scenario recommender systems. Extensive experiments on three public datasets validate that LLM4MSR is not only an effective paradigm compatible with various multi-scenario backbone models, but it is also efficient without fine-tuning LLM and enables real-time recommendation on industrial recommender systems. Besides, the interpretability of conventional recommender system is improved.
Appendix A EXPERIMENTAL SETTINGS
Table 4. The hyper-parameter settings of experiments.
Dataset KuaiSAR-small KuaiSAR Amazon
Embedding dimension 16 16 16
Batch size 4096 4096 1024
Hidden layers of backbone [256, 128, 64] [256, 128, 64] [64, 32, 16]
Learning rate 2e-4 2e-4 2e-3
The hyper-parameters are summarized in Table 4. Meanwhile, the designed template of scenario- and user-level prompt are shown in the purple and blue box in Figure 2, respectively. Specifically, the most recent 55, 55, and 35 positive interactions, i.e., with click, forward, and like behavior on each scenario are leveraged when constructing the user-level prompt on KuaiSAR-small, KuaiSAR, and Amazon dataset because the maximum context length of ChatGLM2-6B is 8K. Besides, the Adam (Kingma and Ba, 2014) optimizer is adopted. The meta network employed is simply a fully connected network. Overall, all the experimental results shown are averaged over 3 runs.
Appendix B Prompt template
In this section, we provide the designed scenario-level and user-level prompt template, where the information in square brackets needs to be filled. Specifically, to capture the commonality among scenarios and emphasize distinction for each scenario, the present scenario-level prompts only differ in the instruction: ‘explicitly summarize the distinction of [scenario_i name]’, where [scenario_i name] is the name of each scenario like recommendation or search scenario in KuaiSAR. Meanwhile, other distinctive information of scenario like prior expert knowledge can be provided. That’s the reason why the prompts are constructed separately for each scenario in Algorithm 1. Prompts with the best performance are:
  • Scenario-level Prompt:
    "On platform [dataset name], [platform description],
    suppose there are [specific number] interactions
    in [scenario_1 name], [specific number] interactions
    in [scenario_2 name]... Besides, in these scenarios there
    are [specific number] users [user description] and
    [specific number] items [item description] where
    [specific number] users and [specific number] products
    are overlapped. From the relationship among these
    scenarios and statistics given, explicitly summarize
    the scenario commonality among the three scenarios
    and the distinction of [scenario name]."
  • User-level Prompt:
    "On platform [dataset name], suppose there are
    [scenario_1 name], [scenario_2 name]... In [scenario_1
    name], user_[id] [historical behavior + item_type
    + category]... Besides, in [scenario_2 name], this user
    [historical behavior + item_type + category].
    For this user, considering the interaction frequency
    and item title plus description information, explicitly
    summarize the common interest among the three scenarios
    first, then summarize the distinct interest in
    each scenario."
ACKNOWLEDGEMENT
This research was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund (No. 2024002)), CCF-BaiChuan-Ebtech Foundation Model Fund, and Kuaishou.